The demand and supply of imported lamb shown as below. Find the CONSUMER's surplus from the equilibrium point. *

P = 70 - 2Q

P = 10 + Q

where

P = price of lamb per kilogram (RM)

Q = quantity of lamb (kg)



1
Expert's answer
2021-09-26T20:14:50-0400
"Q_D=35-0.5p"

"Q_=p-10"

"p=30"

"Q=20"

"CS=\frac{70-30}{2}\times20=400"